OpenStudy (anonymous):

-3x^3 -5x^2 -36 divided by x +3 please divide usug any mean and give only answer

OpenStudy (anonymous):

-3x^2 + 4x - 12

OpenStudy (anonymous):

-3x^3 -5x^2 -36 divided by x +3 = [x +3][Ax^2 + Bx +C ] = [x +3][Ax^3 + Bx^2 +Cx +3Ax^2 +3Bx +3C ] [equate coefficients] -3x^3 -5x^2 -36 =[Ax^3 + Bx^2 +Cx +3Ax^2 +3Bx +3C ] By inspection A = -3 -5 = 3A +B = -9 +B so B= 4 0[no single power of x in LHS] = 3B + C = 12+C so C = -12 -36 = 3C [proven] Answer = [-3x^2 + 4x -12 ]